# AdaDetectGPT: Adaptive Detection of LLM-Generated Text with Statistical Guarantees

### v1

@@ E.4 Proof of Theorem 2
Denote $Z_{t}=\widehat{w}\left(\log q_{t}(X_{t}|X_{<t})\right)-\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}(\bullet|X_{<t})}\widehat{w}\left(\log q_{t}(X_{t}|X_{<t})\right)$. Then if $\bm{X}\sim q$, we have $\mathbb{E}\left\{Z_{t}|X_{<t}\right\}=0$ almost surely. Under Assumptions 3, 5 and the boundedness of $\widehat{w}$ (the B-spline basis are bounded), it is easy to verify that $Z_{t}$ satisfies all conditions of Lemma S1. Therefore, according to Lemma S1, we obtain for any $\alpha\in(0,1)$,
@@
$\leq$ $O\left(\frac{\log L}{\sqrt{L}}\right)+O\big((\mathbb{E}\big|V_{L}-1\big|)^{1/3}\big).$
@@
$\mathbb{E}(\text{FNR}_{\widehat{w}})\leq\alpha+O\left(\frac{\log L}{\sqrt{L}}\right)+O\big((\mathbb{E}\big|V_{L}-1\big|)^{1/3}\big).$
@@ E.5 Proof of Theorem 3
$Y_{j}^{(w)}=w\left(\log q_{t}\left(X_{t}\mid X_{<t}\right)\right),$
$\mu_{j}^{(w)}=\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}}w\left(\log q_{t}\left(\widetilde{X}_{t}\mid X_{<t}\right)\right),$
$(\sigma_{j}^{(w)})^{2}=\textup{Var}_{\widetilde{X}_{t}\sim q_{t}}w\left(\log q_{t}\left(\widetilde{X}_{t}\mid X_{<t}\right)\right).$
@@
$\widehat{\Delta}_{n}(\boldsymbol{x})$ $\leq\frac{1}{\sqrt{L}}\left|\sum_{j=1}^{L}y_{j}^{(\widehat{w})}-\mu_{j}^{(\widehat{w})}\right|\left|\sqrt{L^{-1}\sum_{j=1}^{L}\left(\sigma_{j}^{(\widehat{w})}\right)^{2}}-\sqrt{L^{-1}\sum_{j=1}^{L}\left(\sigma_{j}^{(w^{*})}\right)^{2}}\right|^{-1}$ (34a)
$\hskip 20.00003pt+\left\{\frac{1}{L}\sum_{t=1}^{L}\left(\sigma_{j}^{(w^{*})}\right)^{2}\right\}^{-\frac{1}{2}}\frac{1}{\sqrt{L}}\left|\sum_{j=1}^{L}\left(Y_{j}^{(\widehat{w})}-Y_{j}^{(w^{*})}\right)-\left(\mu_{j}^{(\widehat{w})}-\mu_{j}^{(w^{*})}\right)\right|,$ (34b)
@@
Appendix F Experiment details
Name Model Scale (Billion)
GPT-2 (Radford et al., 2019) openai-community/gpt2-xl 1.5B
GPT-Neo (Black et al., 2021) EleutherAI/gpt-neo-2.7B 2.7B
OPT-2.7 (Zhang et al., 2022) facebook/opt-2.7b 2.7B
GPT-J (Wang & Komatsuzaki, 2021) EleutherAI/gpt-j-6B 6B
Qwen2.5 (Yang et al., 2024a) Qwen/Qwen2.5-7B 7B
Mistral (Jiang et al., 2023) mistralai/Mistral-7B-v0.3 7B
Llama (AIMeta, 2024) meta-llama/Meta-Llama-3-8B 8B
GPT-NeoX (Black et al., 2022) EleutherAI/gpt-neox-20b 20B
{'role': 'system', 'content': 'You are a Review writer on Yelp.'},
and
respectively.
Appendix G Additional numerical results
G.1 Results on additional open-source models
@@
G.2 Computational analysis
From Table S8, the runtime of AdaDetectGPT is around 44 seconds and changes marginally with respect to $d$. This is because we can use a closed-form expression to learn a witness function. This time is nearly negligible compared with the time required to compute logits, which involves feeding tokens from multiple passages into LLMs. Furthermore, the training time when $n$ increases is shown in Table S9. It shows AdaDetectGPT saves more runtime and memory than ImBD.
@@ G.2 Computational analysis
Table S8: Runtime scale with $d$.
@@
ImBD 45.65(9.56) 68.20(9.57) 93.00(9.57) 116.08(9.56) 136.53(9.56) 158.15(9.58)
AdaDetectGPT 9.28(0.36) 23.45(0.37) 40.19(0.37) 44.25(0.37) 59.57(0.60) 69.56(0.37)
Table S9: Runtime (memory in parenthesis) scale with $n$ comparing with ImBD. The runtime is measured in seconds and the memory is measured in GB.
G.3 Analysis factors in training $w$
Since AdaDetectGPT requires training a witness function, we systematically examine three key factors influencing its performance: (1) the size of the training set; (ii) tuning parameters for generating B-spline basis and (iii) distribution shift between training and test data.
Robust performance on various training data sizes. We evaluate AdaDetectGPT across varying dataset sizes by setting $n_{1}=n_{2}\in\{100,200,300,400,500,600\}$ for both human- and machine-generated texts. Figure S5 demonstrates that AdaDetectGPT has a clear performance advantage over FastDetectGPT when sample size is large. This is expected because a larger sample size leads to a more accurate estimation of $w$. Notably, even with limited data $n_{1}=n_{2}=100$, AdaDetectGPT maintains superior accuracy compared to baseline methods, though the performance gap decreases. These results highlight our method’s effectiveness on learning the witness function.
Insensitivity on tuning parameters. The two critical tuning parameters of B-splines are: (i) the number of basis functions (n_base) and (ii) the maximum polynomial order. Our experiments fix one parameter while varying the other (with n_base=16 or order=2 as defaults). As shown in Figure S6 in Appendix, AdaDetectGPT achieves the highest AUC scores so long as n_base $\geq 4$. Besides, enlarging n_base improves the AUC of AdaDetectGPT although the improvement becomes marginal when n_base $\geq 16$. Figure S6 also shows that increasing the polynomial order from linear to quadratic visibly improve the performance; while increasing order from quadratic to cubic/quartic has a limited gain. Having said that, even when the B-splines function is piecewise linear, our method still outperform all baselines.
Robust against distribution shift. In this part, we create training sets with different distributions than the test data by varying the number of human prompt tokens in machine-generated text. In contrary, for the test data, the number of human prompt tokens are fixed. As shown in Figure S7, AdaDetectGPT demonstrates high robustness - the distributional discrepancy between training and test data has negligible effect on classification accuracy. Notably, our method always achieves the highest AUC across all experimental setup.
@@ G.3 Analysis factors in training $w$
G.4 Robustness on adversarial Attacks
We have conducted additional experiments to evaluate the robustness of AdaDetectGPT against 2 adversarial attacks: (i) paraphrasing, where an LLM is instructed to rephrase human-written text, and (ii) decoherence, where the coherence LLM-generated text is intentionally reduced to avoid detection. These experiments were carried out across 3 datasets and 3 types of sampling and scoring models setup, resulting in a total of 18 settings. Both adversarial attacks were implemented following Bao et al. (2024).
paraphrasing decoherence
DetectGPT Xsum Writing PubMed Avg. Xsum Writing PubMed Avg.
Table S10: Detection of LLM texts under two adversarial attacks.
Results are reported in the Table S10. It can be seen that, AdaDetectGPT consistently achieves higher AUCs than Fast-DetectGPT in most of the 18 scenarios. And the improvement reaches up to 10% for paraphrasing and up to 85% for decoherence. These results suggest that AdaDetectGPT attains more robust performance against adversarial attacks.
G.5 Accuracy on open-source models under black-box setting
@@
Appendix H Broader impact and limitation
AdaDetectGPT is a statistically efficient detector for machine-generated text. By accurately detecting generated passages, it can help safeguard AI systems against fake news, disinformation, and academic plagiarism.
In this paper, we prove rigorous guarantees in the white-box setting and demonstrate empirical performance in the black-box setting. The promising results under black-box setting motivate future work to establish matching theoretical guarantees. Moreover, even in the white-box setting, practical LLM’s text generation often employs sampling parameters (e.g. temperature and top_k) that makes the sampling distribution differs from the conditional distribution derived from the source model. This mismatch causes MCLT may not hold in practice. Fortunately, we observe that our test statistic still shifts toward a positive mean (see Figure S8), implying FNR remains control.
Figure S8: Histogram of statistics in three dataset. Each panel visualizes the histogram of statistics. The yellow histogram corresponds to the case when the sampled texts exactly follow the conditional probability of the source model, while purple histogram corresponds to text drawn with deviations from that distribution.

### v2

@@ E.4 Proof of Theorem 2
Denote $Z_{t}=\widehat{w}\left(\log q_{t}(X_{t}|X_{<t})\right)-\mathbb{E}_{\widetilde{X}_{t}\sim q_{t}(\bullet|X_{<t})}\widehat{w}\left(\log q_{t}(X_{t}|X_{<t})\right)$. Then if $\bm{X}\sim q$, we have $\mathbb{E}\left\{Z_{t}|X_{<t}\right\}=0$ almost surely. Without loss of generality, assume $\widehat{w}$ is bounded. Otherwise, we can define $\widehat{w}=\phi^{\top}\widehat{\beta}/\|\widehat{\beta}\|_{2}$ to make it bounded. Under the lower bound assumption in Assumption 3, it is easy to verify that $Z_{t}$ satisfies all conditions of Lemma S1. Therefore, by invoking Lemma S1, we obtain for any $\alpha\in(0,1)$,
@@
$=$ $O\left(\frac{\log L}{\sqrt{L}}\right)+O\big((\mathbb{E}\big|V_{L}-1\big|)^{1/3}\big).$
@@
$\mathbb{E}(\text{FNR}_{\widehat{w}})=\alpha+O\left(\frac{\log L}{\sqrt{L}}\right)+O\big((\mathbb{E}\big|V_{L}-1\big|)^{1/3}\big).$
@@ E.5 Proof of Theorem 3
$Y_{j}^{(w)}=w\left(\log q_{j}\left(X_{j}\mid X_{<j}\right)\right),$
$\mu_{j}^{(w)}=\mathbb{E}_{\widetilde{X}_{j}\sim q_{j}}w\left(\log q_{j}\left(\widetilde{X}_{j}\mid X_{<j}\right)\right),$
$(\sigma_{j}^{(w)})^{2}=\textup{Var}_{\widetilde{X}_{j}\sim q_{j}}w\left(\log q_{j}\left(\widetilde{X}_{j}\mid X_{<j}\right)\right).$
@@
$\widehat{\Delta}_{n}(\boldsymbol{x})$ $\leq\frac{1}{\sqrt{L}}\left|\sum_{j=1}^{L}y_{j}^{(\widehat{w})}-\mu_{j}^{(\widehat{w})}\right|\left|\left[L^{-1}\sum_{j=1}^{L}\left(\sigma_{j}^{(\widehat{w})}\right)^{2}\right]^{-1/2}-\left[L^{-1}\sum_{j=1}^{L}\left(\sigma_{j}^{(w^{*})}\right)^{2}\right]^{-1/2}\right|$ (36a)
$\hskip 20.00003pt+\left\{\frac{1}{L}\sum_{j=1}^{L}\left(\sigma_{j}^{(w^{*})}\right)^{2}\right\}^{-\frac{1}{2}}\frac{1}{\sqrt{L}}\left|\sum_{j=1}^{L}\left(Y_{j}^{(\widehat{w})}-Y_{j}^{(w^{*})}\right)-\left(\mu_{j}^{(\widehat{w})}-\mu_{j}^{(w^{*})}\right)\right|,$ (36b)
@@
E.6 Smooth witness functions
An important quality of spline estimators, which in part motivated our choice of estimator for the witness function, is their ability to learn smooth regression functions at optimal rates. Following the proof of Theorem 2.1 in Chen & Christensen (2015) one can show that when the optimal witness (that is, the function minimizing the functional (21) among all functions with bounded $\ell_{2}$ norm) is $\beta$-Hölder smooth, the estimated witness function attains the rate
$\sup_{z}\left|\widehat{w}(z)-w^{*}(z)\right|=O_{P}\left(\sqrt{\frac{d\log(n)}{n}}\right)+O(d^{-\beta}).$
Consequently, choosing the number of spline bases as $d=\Theta((n/\log(n))^{\frac{1}{2\beta+1}})$ the sup norm loss will be of the order $O_{P}\left((\log(n)/n)^{\frac{\beta}{2\beta+1}}\right)$. This rate was shown to be optimal by Stone (1982). One can therefore show that choosing the number of spline bases in this way the expected true negative rate will be lower bounded as
$\mathbb{E}(\text{TNR}_{\widehat{w}})\geq\text{TNR}_{w^{*}}-O\left((\log(n)/n)^{\frac{\beta}{2\beta+1}}\right).$
In this section we provide a sketch of this result.
We argue along the same lines as the proof of Theorem 3. Since $\mathbb{E}(\text{TNR}_{\widehat{w}})\geq\text{TNR}_{w^{*}}-\mathbb{E}(|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}|)$ is enough to upper bound $\mathbb{E}(|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}|)$. Define the quantity $\widehat{\Delta}_{n}=\sup_{\boldsymbol{x}}|\widehat{T}_{n}(\boldsymbol{x})-T^{*}(\boldsymbol{x})|$, and introduce the event
$A(\kappa_{1})=\left\{\widehat{\Delta}_{n}\leq\kappa_{1}\times(d^{-\beta}+\left(\log(n)/n\right)^{\frac{\beta}{2\beta+1}})\right\}$ (40)
Therefore, with arbitrary $\kappa_{1}$ for some absolute $\kappa_{2}$ we have that
$\mathbb{E}(|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}|)$ $=\mathbb{E}(|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}|\times\mathbf{1}_{\{A(\kappa_{1})\}})+\mathbb{E}(|\text{TNR}_{\widehat{w}}-\text{TNR}_{w^{*}}|\times\mathbf{1}_{\{\neg A(\kappa_{1})\}})$
$\leq 2\mathbb{E}(\mathbb{P}_{\boldsymbol{X}\sim p}(|T^{*}(\boldsymbol{X})-z_{\alpha}|\leq\widehat{\Delta}_{n})\times\mathbf{1}_{\{A(\kappa_{1})\}})+\mathbb{P}(\neg A(\kappa_{1}))$ (41a)
$\leq\kappa_{1}\mathbb{E}(\hat{\Delta}_{n}\mid A(\kappa_{1}))+\mathbb{P}(\neg A(\kappa_{1}))$ (41b)
where (41a) follows from the arguments leading up to (34) and (41b) holds on the event (35). Using the mean value theorem and Assumption 3 one can show that up to constants $\hat{\Delta}_{n}$ is smaller than $\sup_{z}\left|\widehat{w}(z)-w^{*}(z)\right|$, therefore by (40) the first term in (41b) will be of the order $O(d^{-\beta}+\left(\log(n)/n\right)^{\frac{\beta}{2\beta+1}})$. Examining the proof of Theorem 2.1 in Chen & Christensen (2015) it can be seen that for any $\kappa_{3}>0$ on may choose $\kappa_{1}$ so that $\mathbb{P}(\neg A(\kappa_{1}))<n^{-\kappa_{3}}$. Therefore, we can choose $\kappa_{1}$ appropriately so that the second term in (41b) will be of the order $o(\left(\log(n)/n\right)^{\frac{\beta}{2\beta+1}})$. Finally letting $d=\Theta((n/\log(n))^{\frac{1}{2\beta+1}})$ yields the desired result.
As mentioned in the main text, when the optimal witness function is believed to be smooth but the order of the smoothness is not known the number of spline bases can be chosen in a data driven way via Lepski’s method (Lepski & Spokoiny, 1997).
Appendix F Additional numerical results
F.1 Results on additional open-source models
@@
F.2 Additional results on black-box setting
Table S8: AUC scores of various detectors to detect text generated by Gemini-2.5 across datasets.
Method XSum Writing Yelp Essay Avg. RoBERTaBase 0.5311 0.5202 0.5624 0.7279 0.5854 RoBERTaLarge 0.6583 0.5888 0.6029 0.8180 0.6845 Likelihood 0.7127 0.7547 0.6566 0.7565 0.7201 Entropy 0.5754 0.5088 0.6023 0.6038 0.5726 LogRank 0.6084 0.5743 0.6896 0.7504 0.6557 LRR 0.5960 0.5382 0.5580 0.6703 0.5906 Binoculars 0.8500 0.9453 0.9698 0.9908 0.9390 RADAR 0.8184 0.5152 0.6300 0.5891 0.6382 BiScope 0.7633 0.6800 0.7097 0.8167 0.7642 Fast-DetectGPT 0.8404 0.9443 0.9695 0.9914 0.9364 AdaDetectGPT 0.8432 0.9484 0.9644 0.9947 0.9377 Relative ( ) 1.7544 7.4163 37.8238 1.9916
F.3 Computational analysis
In this part, we study the runtime for learning the witness function. From Table S9, the runtime of AdaDetectGPT is around 44 seconds and changes marginally with respect to $d$. This is because we can use a closed-form expression to learn a witness function. This time is nearly negligible compared with the time required to compute logits, which involves feeding tokens from multiple passages into LLMs. Furthermore, the training time when $n$ increases is shown in Table S10, and we can see that the runtime for training is typically no more than one minute.
Table S9: Runtime scale with $d$.
@@ F.3 Computational analysis
Table S10: Runtime (memory in parenthesis) scale with $n$. The runtime is measured in seconds and the memory is measured in GB.
@@
Runtime (Memory) 9.28(0.36) 23.45(0.37) 40.19(0.37) 44.25(0.37) 59.57(0.60) 69.56(0.37)
F.4 Sensitivity analysis
Since AdaDetectGPT requires training a witness function, we examine three factors influencing its performance: (1) the size of the training set; (ii) tuning parameters for generating B-spline basis and (iii) distribution shift between training and test data.
Robust to training data sizes. We evaluate AdaDetectGPT across varying dataset sizes by setting $n_{1}=n_{2}\in\{100,200,300,400,500,600\}$ for both human- and machine-generated texts. Figure S5 demonstrates that AdaDetectGPT clearly outperforms FastDetectGPT when sample size is large. This is expected because a larger sample size leads to a more accurate estimation of $w$. Notably, even with limited data $n_{1}=n_{2}=100$, AdaDetectGPT maintains superior accuracy compared to baseline methods, though the performance gap decreases. These results highlight our method’s effectiveness on learning the witness function.
Insensitivity to tuning parameters. B-spline relies on two critical tuning parameters: (i) the number of basis functions (n_base) and (ii) the maximum polynomial order. Our experiments fix one parameter while varying the other (with n_base=16 or order=2 as defaults). As shown in Figure S6 in Appendix, AdaDetectGPT achieves the highest AUC scores so long as n_base $\geq 4$. Besides, enlarging n_base improves the AUC of AdaDetectGPT although the improvement becomes marginal when n_base $\geq 16$. Figure S6 also shows that increasing the polynomial order from linear to quadratic visibly improve the performance; while increasing order from quadratic to cubic/quartic has a limited gain. Finally, even when the B-spline basis is set to a piecewise linear function, our method still outperform all baselines.
Robust against distribution shift. We create training datasets with different distributions from the test data by varying the number of human prompt tokens in machine-generated text. In contrast, for the test data, the number of human prompt tokens are fixed. As shown in Figure S7, AdaDetectGPT demonstrates high robustness to the distributional discrepancy between training and test data. It achieves the highest AUC across all experimental setup.
@@ F.4 Sensitivity analysis
F.5 Detecting open-source models in black-box setting
@@
Appendix G Broader impact and limitation
AdaDetectGPT is a computationally and statistically efficient detector for machine-generated text, thus safeguarding AI systems against fake news, disinformation, and academic plagiarism.
Despite AdaDetectGPT’s strong empirical performance in the black-box setting, its theoretical guarantees are mainly established in the white-box setting. Even when restricting to the white-box setting, LLM text generation often involves sampling parameters (e.g., temperature and top_k). Using different parameter values can cause the sampling distribution to deviate from that of the target model we aim to detect. This mismatch invalidates MCLT in practice. Fortunately, we observe that the shape of our statistic remains similar, but shifts toward a positive mean (see Figure S8), implying that FNR control under MCLT remains valid, although being more conservative.
Figure S8: Histogram of our statistics in three datasets. Each panel visualizes the histogram in one dataset. The yellow histogram corresponds to the case when the sampled texts exactly follow the conditional probability of the source model, while purple histogram corresponds to text drawn with from a distribution with different sampling temperatures.